# Strong geometry: knots

### v1

@@ 4.2. Main result
(B) The rotation order of the strands around a multiple intersection and
(C) the order of the intersections (if more than one) in a given arc.
For (A), let $\beta=\pi_{x_{0}}([x_{i},x_{i+1}])$ and $\beta^{\prime}=\pi_{x_{0}}([x_{i^{\prime}},x_{i^{\prime}+1}])$ be two arcs in the line diagram. Suppose that these arcs intersect and that this intersection is simple. Then, $\beta$ is over $\beta^{\prime}$ if the Radon partition of the set $\{x_{0},x_{i},x_{i+1},x_{i^{\prime}},x_{i^{\prime}+1}\}$ is given by $\{x_{0},x_{i},x_{i+1}\}\sqcup\{x_{i^{\prime}},x_{i^{\prime}+1}\}$, see Figure 23.
Figure 23. Radon partition and crossing from the witness.
Furthermore, if $\pi_{x_{0}}([x_{i},x_{i+1}])$ and $\pi_{x_{0}}([x_{i^{\prime}},x_{i^{\prime}+1}])$ intersect, the oriented arcs (from $x_{i}$ to $x_{i+1}$ and from $x_{i^{\prime}}$ to $x_{i^{\prime}+1}$) on the sphere meet at a positive crossing if and only if $\chi(x_{i},x_{i+1},x_{0},x_{i^{\prime}+1})=+1$. For a multiple intersection, the Radon partition determine which arc is above/below the others. We can thus record the order of the new intersections (and on which arc) after the local modfications explained above.
For (B), we notice that the rotation order of several strands around a multiple point only depends on the set of signs of the pairs of crossings which can be determined by the same argument as in (A).
For (C), suppose that both arcs $\beta^{\prime}=\pi_{x_{0}}([x_{i^{\prime}},x_{i^{\prime}+1}])$ and $\beta^{\prime\prime}=\pi_{x_{0}}([x_{i^{\prime\prime}},x_{i^{\prime\prime}+1}])$ intersect arc $\beta=\pi_{x_{0}}([x_{i},x_{i+1}])$. The order of the intersection, say from $\pi_{x_{0}}(x_{i})$ to $\pi_{x_{0}}(x_{i+1})$ is given by the chirotope $\chi_{x_{0}}$ according with the rule given in Figure 24. We notice that, by Proposition 2, $\chi_{x_{0}}$ can be completely determined from $\chi$ and $\chi_{\Lambda}$.
@@ 5. Further research directions
Ringel’s isotopy conjecture asked whether any two arrangements of lines with the same topological type can be isotoped into each other. In other words, Ringel’s conjecture asked whether the realization space of rank 3 matroids is connected. The famous Mnëv Universality Theorem [8] provides a spectacular negative solution, it states that every semi-algebraic variety is stably equivalent to the realization space of some oriented matroid, that is, the realization space of oriented matroids can have the topology of any semi-algebraic set and, in particular, it can be disconnected. As mentioned above, oriented matroids do not capture the combinatorics of the cell configuration induced by the spanned lines contrary to strong geometries that completely does. In the same flavor as for oriented matroids, we investigate [5] a universality result à la Mnëv for rank 3 strong geometries.
The spherical line diagrams were introduced to deal with the fact that the radial projection is from a point belonging to the polygonal knot. This might be avoided if given two set of points $P$ and $Q$ having the same strong geometry then one can show the existence of two points $p$ and $q$ such that $P\cup\{p\}$ and $Q\cup\{q\}$ also have the same strong geometry. In this case, we just can add this additional point from which the radial projection is applied. Figure out the existence of these new points is not difficult when $d=2$ but it is not straightforward for $d\geq 3$.
It turns out that the notion of strong geometry can be generalized in the planar case as follows. We consider the lines spanned by both the set of the original points and the points arising from line intersections. The latter induce a new set of lines and therefore new intersection points from which new lines can be spanned and so on. These generalizations happen to be closely related to the notion of $k$-equivalence between two configurations of points (the case $k=1$ concerns the existence of a new point to be added to a strong geometry, as discussed above). A notion of $\infty$-equivalence also naturally emanate.
Finally, as pointed in the proof of Claim 1, we introduce a new way to show that a knot diagram can be determined by the corresponding Gauss diagram. We believe that this approach can be extended to investigate spatial graphs and virtual knots.

### v2

@@ 4.3. Main result
(B) the order of the intersections (if more than one) along a given arc in $\mathcal{K}$.
For (A), let $\beta=\pi_{x_{0}}([x_{i},x_{i+1}])$ and $\beta^{\prime}=\pi_{x_{0}}([x_{i^{\prime}},x_{i^{\prime}+1}])$ be two arcs in $\mathcal{K}$. Suppose that these arcs intersect in a simple point. Then, $\beta$ is over $\beta^{\prime}$ if the Radon partition of the set $\{x_{0},x_{i},x_{i+1},x_{i^{\prime}},x_{i^{\prime}+1}\}$ is given by $\{x_{0},x_{i},x_{i+1}\}\sqcup\{x_{i^{\prime}},x_{i^{\prime}+1}\}$, see Figure 22.
Figure 22. Radon partition and crossing in $\mathcal{K}$ from the witness $x_{0}$.
Furthermore, if arc $\pi_{x_{0}}([x_{i},x_{i+1}])$ intersect arc $\pi_{x_{0}}([x_{i^{\prime}},x_{i^{\prime}+1}])$ then the oriented arcs (from $x_{i}$ to $x_{i+1}$ and from $x_{i^{\prime}}$ to $x_{i^{\prime}+1}$) on the sphere meet at a positive crossing if and only if $\chi(x_{i},x_{i+1},x_{0},x_{i^{\prime}+1})=+1$.
For (B), suppose that both arcs $\beta^{\prime}=\pi_{x_{0}}([x_{i^{\prime}},x_{i^{\prime}+1}])$ and $\beta^{\prime\prime}=\pi_{x_{0}}([x_{i^{\prime\prime}},x_{i^{\prime\prime}+1}])$ intersect arc $\beta=\pi_{x_{0}}([x_{i},x_{i+1}])$. The order of the intersection, say from $\pi_{x_{0}}(x_{i})$ to $\pi_{x_{0}}(x_{i+1})$ is given by the chirotope $\chi_{x_{0}}$ according with the rule given in Figure 23. We notice that, by Proposition 3, $\chi_{\wedge,x_{0}}$ can be completely determined from $\chi_{\Lambda}$.
@@
5. Spatial graphs
A spatial graph is an embedding of a finite graph in $\mathbb{R}^{3}$, that is, the vertices are distinct points and the edges are simple curves between them in such a way that any two curves are either disjoint or meet at a common vertex. A spatial graph is called linear if each edge is a straight line segment. Two spatial graphs are said to be equivalent if they are ambiently isotopic [17, 21].
Theorem 2.
Let ${R}(G)$ and ${R}^{\prime}(G)$ be two linear spatial representations of $G$. Let $X$ and $X^{\prime}$ be the corresponding set of points in $\mathbb{R}^{3}$ associated to the vertices in such representations. If $\mathsf{SGeom}_{\mathsf{Aff}}(X)$ and $\mathsf{SGeom}_{\mathsf{Aff}}(X^{\prime})$ are isomorphic then ${R}(G)$ is equivalent to ${R}^{\prime}(G)$.
Let us briefly recall the notion of spatial graphoids needed for the proof of Theorem 2.
A graphoid is an equivalence class of graphoid diagrams. A graphoid diagram is a graph with various distinguished degree-one vertices generically immersed in $\mathbb{S}^{2}$, where each double point is decorated as a classical crossing labeled over/under, see Figure 24.
Figure 24. A graphoid with four distinguished degree-one vertices.
Graphoids have been studied in different contexts, for instance, in connection with a topological structure that arises in proteins [15]. A Gauss code is associated to any graphoid diagram $\mathcal{G}$ as follows. We choose an orientation for each edge. We then write down the Gauss code for the shadow by recording whether each arrival at a crossing is an over-crossing (O) or an under-crossing (U), we can also label each crossing by its sign $+$ or -- according to handedness’ rule, see Table 1. The difference with usual knots is that we add one more piece of information. For each vertex $v$ of $G$, we consider the sequence of edges incident to $v$, appearing in counterclockwise order around $v$. These edges sequences are considered up to cyclic permutation.
Table 1. Gauss code of the spatial graphoid given in Figure 24.
Gauss code
$a1AO_{+}b$
$a2BU_{-}FO_{+}EU_{+}i$
$a3BO_{-}AU_{+}d$
$b4FU_{+}h$
$b5c$
$b6DU_{-}CO_{-}e$
$d7DO_{-}CU_{-}EO_{+}h$
$d8i$
$i9h$
$h10g$
$h11f$
More conveniently, Gauss codes can be understood as Gauss ‘arrow’  diagrams (as for knotoids). We also consider edges sequences around each vertex. We label points along the edges according to the crossing sequences of the Gauss code and draw an arrow between each pair of occurrences of the same label. The arrow is oriented from the undercrossing edge to the overcrossing edge and labeled with the sign of the crossing. We notice that once the arrows are drawn, the labels on the edges are redundant, and can thus be removed. In general, there is no canonical choice for ‘the immersion’of the graph. Even if the underlying graph has a planar embedding, the order of the edges around the vertices may prevent us from using it, see Table 2.
Table 2. Gauss diagram of the graphoid diagram given in Figure 24.
Edge sequences
$a:1,3,2$
$b:1,4,6,5$
$c:5$
$d:3,7,8$
$e:6$
$f:11$
$g:10$
$h:4,11,10,9,7$
$i:8,2,9$
Lemma 4.
A graphoid diagram is determined by its Gauss diagram.
Proof.
We essentially follow the same arguments as those given in the proof of Lemma 3. We only need to adapt the construction of a valid travel $T$ in the Gauss diagram of a spatial graphoid. We have to determine the rule when $T$ gets to a vertex of degree 3$\geq$ 3≥ 3 (not existing in knotoids), that is, we have to determine the turns of the corresponding walk $W$ when going through vertices in $\mathcal{G}$. For this, we use the edge sequence information, if $W$ arrives at $v$ along an edge $e$ then it leaves $v$ by following the edge $f$ appearing before $e$ in counterclockwise direction, that is, the next edge in clockwise direction. This new rule forces that $W$ always turns «left» at each vertex staying in the same face, see Figure 25.
Figure 25. Turn of $W$ at vertex $v$.
Remark 3.
Lemma 4 can be naturally extended to virtual graphoids, that is, graphoids having virtual crossings. The virtual notion was introduced by Kauffman in his seminal paper [16]. The idea is not that a virtual crossing is just an ordinary degree-four vertex but rather that a virtual crossing is not really there (hence the name ‘virtual’).
We may now prove Theorem 2.
Proof of Theorem 2. We shall mimic the proof of Theorem 1. Let $X=(x_{0},\dots,x_{n-1})$ be the tuple of points in $\mathbb{R}^{3}$ associated to the vertices $(v_{0},\dots,v_{n-1})$ in the spatial representation of $R(G)$. Let $\pi_{x_{0}}(R(G))$ be the radial projection of $R(G)$ emitted from $x_{0}$. It can be checked that such projection induces a graphoid, say $\mathcal{G}$, with as many distinguished vertices as the degree of the vertex corresponding to point $x_{0}$ (this can be justified in a similar way as for knotoids arising from radial projections of polygonal knots). If the diagram is not connected, then the corresponding components are unlinked, and a sequence of isotopies for each component leads to an isotopy for the whole spatial graph. Therefore, we can suppose without loss of generality that the diagram is connected. This implies that the faces of the diagram have connected boundaries and thus the arguments used in the proof of Theorem 1 work similarly in this case.
Let $\mathcal{G}^{\prime}$ be the spatial graphoid arising from $R^{\prime}(G)$. We notice that any planar isotopy between the spatial graphoids $\mathcal{G}$ and $\mathcal{G}^{\prime}$ can be extended to an isotopy between $R(G)$ and $R^{\prime}(G)$ by fixing $x_{0}$ and keeping the segments $[x_{0},x_{i}]$, for each $x_{i}$ where its associated vertex $v_{i}$ is a neighbor of $v_{0}$, straight while doing the isotopy. Therefore, by Lemma 4, it is enough to show that the Gauss diagram associated to $\mathcal{G}$ is completely determined by $\mathsf{SGeom}_{\mathsf{Aff}}(X)=(M_{\mathsf{Aff}}(X),M_{\mathsf{Wedge}}(\mathsf{in}(X)))$ (implying that $R(G)$ can also be uniquely reconstructed from such strong geometry).
As discussed in the proof of Theorem 1, $\mathsf{SGeom}(X)_{\mathsf{Aff}}$ determines in the diagram : the pairs of intersecting arcs (and which one is over/under the other one) together with the corresponding sign, the local type of multiple intersection and thus their local modifications and the order of the intersections (if more than one) along a given arc. Although all of these are needed to determine $\mathcal{G}$, we still have to show that $\mathsf{SGeom}(X)_{\mathsf{Aff}}$ also determines the order of appearance, say counter clockwise, of the edges around each vertex. In fact this order is determined by $M_{\mathsf{Aff}}(X)$ itself. Indeed, the two pieces of oriented equators $\pi_{x_{0}}([x_{i},x_{j}])$, $\pi_{x_{0}}([x_{i},x_{j^{\prime}}])$ meet either positively or negatively depending on the sign of $\Delta(\pi_{x_{0}}(x_{i}),\pi_{x_{0}}(x_{j}),\pi_{x_{0}}(x_{j^{\prime}}))=\chi(x_{0},x_{i},x_{j},x_{j^{\prime}})$. But the data of the signs of such crossings between pieces of equators is equivalent to the order at which the $x_{j}$’s appear around $x_{i}$ (on the unit sphere centered on $x_{0}$). This is due to the fact that the cyclic order of the strands is given by the sign of each pair (as explained above) and, by Proposition 1, the sign of a pair $(x_{i}\wedge x_{j}),(x_{i}\wedge x_{j^{\prime}})$ is given by the sign $\chi(x_{i},x_{j},x_{j^{\prime}})$. Essentially, the circular order around $x_{i}$ is the ‘twofold witnessed’ (first by $x_{0}$ and then by $x_{i}$) rank 222 chirotope. $\square$
We end with the following
Question 2.
Let $R(G)$ be a linear spatial representations of $G$. Let $X$ be the set of points in $\mathbb{R}^{3}$ associated to the vertices of such representation. Is it true that $R(G)$ is determined by $M_{\mathsf{Aff}}(X)$
@@ 6. Further research directions
6.1. Universality
Ringel’s isotopy conjecture asked whether any two arrangements of lines with the same topological type can be isotoped into each other. In other words, Ringel’s conjecture asked whether the realization space of rank 3 matroids is connected. The famous Mnëv’s Universality Theorem [19] provides a spectacular negative solution, it states that every semi-algebraic variety is stably equivalent to the realization space of some oriented matroid, that is, the realization space of oriented matroids can have the topology of any semi-algebraic set and, in particular, it can be disconnected. As mentioned above, oriented matroids do not capture the combinatorics of the cell configuration induced by the spanned lines contrary to strong geometries that completely does. In the same flavor as for oriented matroids, an ‘inclusion-type’ generalization of the universality result à la Mnëv for rank 3 strong geometries is investigated [14].
6.2. $k$-equivalence
The notion of knotoid is needed to deal with the fact that the radial projection is from a point belonging to the polygonal knot. This might be avoided if given two set of points $P$ and $Q$ having the same strong geometry then one can show the existence of two points $p$ and $q$ such that $P\cup\{p\}$ and $Q\cup\{q\}$ also have the same strong geometry. In this case, we just can add this additional point from which the radial projection is applied. Figuring out the existence of these new points is not difficult when $d=2$ but it is not straightforward for $d\geq 3$.
It turns out that the notion of strong geometry can be generalized in the planar case as follows. We consider the lines spanned by both the set of the original points and the points arising from line intersections. The latter induce a new set of lines and therefore new intersection points from which new lines can be spanned and so on. These generalizations happen to be closely related to the notion of $k$-equivalence between two configurations of points. A notion of $\infty$-equivalence also naturally emanates. Such equivalences are studied in [12] in connection with some asymptotic rigidity properties. The case $k=2$ is related to the existence of a new point to be added to a strong geometry, as discussed above.
@@ 6.2. $k$-equivalence
6.3. Adjoints
Figure 1 gives an exemple of an oriented matroid of rank 3 with two realizations that generate non-isomorphic adjoints. Moreover, there exists examples of two non-isomorphic simple oriented matroids of rank 3 that have isomorphic adjoints, see [1, 2] and [4, Exercice 5.13 (a)]. In the realizable case, Corollary 2 asserts that if the adjoints are isomorphic then the oriented matroids are equal or opposite to each other.